# How to build the best medical image segmentation algorithm using foundation models: a comprehensive empirical study with Segment Anything Model

### v2

@@ Abstract
1 Introduction
@@
Medical image segmentation refers to identifying the region of interest, typically anatomical structures or tumors, in an image. This process is essential in the field of medical image analysis, contributing significantly to the effectiveness of computer-aided diagnosis [1]. The introduction of deep learning techniques has led to a variety of automated segmentation methods [57, 85], further enhancing the precision and efficiency of analyzing medical images. Despite the progress made, it is still difficult to determine the most effective method for developing an optimal segmentation model, especially with the recent introduction of vision foundation models dedicated to segmentation. Therefore, in our study, we aim to explore and define the optimal strategies for developing automated segmentation algorithms specifically designed for medical images and with a focus on foundation models.
@@
In the era of deep learning, a common practice for automated medical image segmentation is to directly train UNet [70] or its variants [75, 83] for a dataset and task. Recently, there has been a growing interest in the field of customizing vision foundation models towards specific segmentation tasks [96, 29]. Foundation models are (typically transformer-based [82]) neural networks pre-trained with extremely large datasets, such that they can generalize to a wide range of tasks and datasets. The seminal Segment Anything Model (SAM) [45] was proposed as the first large model designed with segmentation tasks in mind. SAM employs a prompt-based training method, aiming to generate target object’s segmentation masks in response to user-provided prompts, e.g. points or bounding boxes. It has demonstrated comparable or even superior zero-shot performance to fully supervised models in processing natural images. However, SAM performance is less impressive when applied directly to medical images [61, 21, 39], and the requirement of prompts makes its direct usage with medical imaging segmentation tasks challenging.
@@
Removing the prompt requirement in SAM is relatively straightforward as we can use “dummy” prompt embeddings as inputs during fine-tuning, a strategy verified by concurrent work [96]. However, the approaches proposed for adapting SAM to medical imaging tasks vary widely. Some works focused on adapting SAM’s decoder only [92], while others chose to modify the entire network [96]. Some works introduced an additional pre-training stage [59] using multiple medical images while others usually do not. Though some prior survey work [49, 98] have attempted to list different strategies and categorize them into various groups, questions remain regarding (1) which fine-tuned strategy is the most effective, (2) whether these models can achieve better performance than traditional UNets and (3) whether using additional data is helpful.
@@
In this work, we aim to answer these questions by considering three common dataset availability scenarios in the medical imaging field: (1) only a single labeled dataset; (2) multiple labeled datasets for different tasks; and (3) multiple labeled and unlabeled datasets. The first scenario reflects the most common deep learning scenario of adapting SAM to the dataset of interest, and the latter two scenarios bring the opportunity to first incorporate broad medical domain knowledge into SAM using supervised/self-supervised training before adapting it to a specific dataset. Figure 1 summarizes the strategies and experiment setups we include. To avoid terminology confusion, we name these strategies as task-specific supervised training, task-expansive prompt-based learning, and task-agnostic self-supervised learning respectively.
@@
When considering the design choice during task-specific supervised training, our work aligns with several concurrent works [49, 98, 99]. However, our work differs from theirs in that we conduct comprehensive experiments over medical imaging datasets across various modalities for a fair comparison, rather than simply compiling disparate results, and provide an organized and new catalog that could summarize existing approaches. The choice of each category is shown in Figure 2 and will be discussed in Section 2.2. Our category definition reflects the common design choices when building a segmentation algorithm from SAM, and we can categorize most existing works based on the proposed criteria, shown in Table 1.
@@
When incorporating extra medical knowledge into SAM, we explore two main setups of pre-training approaches: task-expansive prompt-based learning and task-agnostic self-supervised learning, depending on the availability of labels. Note that our objective is not to propose a novel pre-training strategy, as such strategies are often constrained by the availability of computation resources and datasets. Instead, our focus is on identifying if additional pre-training with more available medical data is helpful and if it is, to determine what is the best combination of pre-training and task-specific fine-tuning. Therefore, we adopt the strategies used during SAM’s development and one of the most commonly applied pre-training strategies in this field. Namely, we use the same interactive training pipeline where prompts are simulated from the masks for task-expansive prompt-based learning; and masked autoencoding (MAE) [31] as the objective for task-agnostic self-supervised learning. We evaluate these approaches on the same 17 datasets.
@@
Our study is extended to two crucial scenarios when deploying a segmentation algorithm for annotation purposes: few-shot learning and interactive segmentation. The first scenario reflects the case of annotating a new dataset with only a limited number of annotations, while the latter examines the effectiveness of using tools to assist the annotation process. In both scenarios, we limit the evaluation datasets to MRI. When conducting few-shot learning, we limit the number of training samples to 5 and re-examine the effectiveness of setups 1-3 described in Figure 1. In the interactive segmentation scenario, we focus on task-specific supervised training with either box or point as input prompts.
@@
In conclusion, we systematically study different approaches to fine-tune SAM in the medical imaging field. Our paper investigates an important question: what is the best practice to develop an automated segmentation algorithm based on foundation models? We find the answers depend on the availability of the datasets and consider three scenarios: (1) only a single labeled dataset; (2) multiple labeled datasets (with different objects of interest); and (3) multiple labeled and unlabeled datasets. We propose task-specific supervised training, task-expansive prompt-based learning followed by task-specific supervised training, and task-agnostic self-supervised learning followed by task-specific supervised training as the solution to each scenario respectively. By conducting systematic experiments on each strategy, we observe several useful guidelines and summarize them in Section 5.
@@ 2.1 SAM and removal of prompts
To begin with, we first briefly present an overview of the Segment Anything Model (SAM) [45] before introducing the fine-tuning strategies. SAM consists of three components: (1) image encoder, (2) prompt encoder, and (3) mask decoder. The image encoder, built on a robust Vision Transformer (ViT) framework, effectively converts a 2D image into a latent feature representation. Notably, in SAM’s original architecture, the image encoder offers flexibility in model size, with options including ViT-H(uge), ViT-L(arge), and ViT-B(ase) [22]. The prompt encoder is flexible, accommodating different prompt types such as point, box, or mask prompts. Depending on the input type, it generates either sparse embeddings for point/box prompts or dense embeddings for mask prompts. Lastly, the mask decoder is a transformer decoder using two-way cross-attention that integrates image and prompt embeddings to produce a multi-channel mask output, each channel representing varying confidence levels.
@@ 2.1.1 Interactive segmentation to automated segmentation
where the “sparse_embeddings” and “dense_embeddings” serve as inputs to SAM’s mask decoder. Empirical evidence shows the effectiveness of this simple approach [96]. More advanced methods, including adding a detection network to automatedally generate box prompts or a self-prompt module into SAM’s structure [87, 27, 72], are feasible but will introduce extra complexity into SAM’s structure, hampering our analysis on the effectiveness of fine-tuning strategies. Moreover, we believe our work offers a potentially effective strategy, i.e. feeding the generated prompts to the fine-tuned SAM instead of the original one, for this research direction.
@@ 2.2.1 Image Encoder Architecture
In this work, we select ViT-B and ViT-H from SAM’s provided checkpoints because they represent the most efficient and effective versions of SAM’s encoder respectively. Additionally, we explore a more compact option: ViT-T(iny). This backbone is introduced in MobileSAM [94], where the authors distill the knowledge from SAM’s pre-trained ViT-H to ViT-T, a customized vision transformer with of the network parameters and offers an inference speed times faster compared to the original SAM. The lightweight nature can widen the application of SAM, ease the computational requirements, and, more importantly, raise interesting research questions, such as whether distillation can maintain or even lead to the superior performance of SAM during fine-tuning.
@@ 2.2.2 Model Component
Some methods state the effectiveness of maintaining the encoder’s weights and only adapting the lightweight mask decoder of SAM’s architecture, as shown in Table 1. These approaches are grounded in the understanding that SAM’s encoder, trained on a comprehensive collection of natural images, is adequately powerful and knowledgeable in the general imaging domain with zero-shot ability to perform feature extraction, and updating the lightweight decoder can enhance performance in specific applications. Also, it is suggested that fine-tuning the encoder of SAM would demand considerably higher computational resources. However, due to the significant appeal difference between natural images and medical images, it remains unclear if SAM’s encoder could effectively capture the distinctive features of medical images without any modifications to the image encoder. Therefore, we consider these two choices: mask decoder only and mask decoder plus image encoder.
@@
| Decoder + Vanilla | SAM for medical images? [39] | COSMOS 1050K | ViT-H / ViT-B | |
| Generalist Vision Foundation [73] | Multiple | ViT-H | ||
| How to efficiently [38] | ACDC [8] | ViT-H | ||
| All-in-SA [19] | Monuseg [47] | ViT-H | ||
| AdaptiveSAM [64] | Multiple | ViT-B | ||
| SkinSAM [37] | HAM10000 [80] | ViT-B | ||
| DeSAM [27] | Multi-Prostate | ViT-H | ||
| SurgicalSAM [92] | EndoVis [3] | ViT-H | ||
| SAM3D [11] | Multiple | ViT-B | ||
| SAM Fewshot [88] | Multiple | N/A | ||
| Decoder + LoRA | BLO-SAM [97] | Multiple | ViT-B | |
| Encoder + Vanilla | AutoSAM [72] | Multiple | ViT-H | |
| SAMUS [56] | Multiple | N/A | ||
| En/Decoder + Vanilla | Polyp-SAM [55] | Multi-Colonoscopy | ViT-L / ViT-B | |
| CellViT [34] | Multi-nuclei | ViT-H | ||
| En/Decoder + LoRA | Customized SAM [96] | Multi-Organ | ViT-H / ViT-B | |
| MA-SAM [14] | BTVC [24] | ViT-H | ||
| Cheap Lunch [25] | BraTS 18 [62] | ViT-B | ||
| En/Decoder + Adapter | Medical SAM Adapter [86] | BTVC [24] | ViT-H | |
| SAM Fails to Segment? [17] | kvasir-SEG [42] | ViT-H | ||
| Learnable Ophthalmology SAM [67] | Multiple | N/A | ||
| Auto-Prompting SAM [53] | Multiple | LightWeight ViT | ||
| SAM-Med2D [18] | SAM-Med2D [18] | ViT-B | ||
| 3DSAM-adapter [28] | Multi-CT | ViT-B | ||
| SAM-Path [95] | Multiple | LightWeight ViT | ||
| Self-Sampling Meta [51] | ABD-30 [48] | ViT-B | ||
| Cross-modality [74] | MRI | ViT-B |
@@ 2.2.3 Vanilla or Parameter-efficient Fine-tuning Method
When conducting model fine-tuning or transfer learning, the easiest and most intuitive way is to update all model parameters involved when feeding medical domains of data, which is vanilla fine-tuning [6].
Recently, parameter-efficient fine-tuning (PEFT) techniques have gained an increasing amount of attention. Compared with vanilla fine-tuning, PEFT keeps most of the network parameters fixed and only updates a small percentage, i.e. less than , of them. Such a strategy significantly reduces the computational cost and is beneficial in preventing catastrophic forgetting as well as over-fitting [23]. The simplest PEFT techniques involve selecting and fine-tuning a subset of the parameters from the original pre-trained model. These selective methods concentrate on identifying the optimal parameters for fine-tuning, aiming to strike an ideal balance between parameter efficiency and overall model performance. For instance, BitFit [93] proposed fine-tuning only the bias term in a transformer-based network, and Tinytl [12] shares a similar idea but focuses on CNN-based networks. Touvron et al. [79] have demonstrated the effectiveness of selectively fine-tuning only the attention layers in Vision Transformers (ViTs), whereas alternative approaches suggest that exclusively optimizing the normalization layers [7].
Rather than updating certain selected layers, recent research [54] indicates that incorporating extra modules into the large models appears to be more beneficial because it re-configures models towards a downstream task than solely updating network parameters. Among these additional module methods, Adapters, which were originally introduced for learning multi-domain representations [68, 69], have emerged as a particularly prevalent choice. In addition to the conventional Adapter block [35], Low-Rank Adaptation (LoRA), known for its efficient low-rank matrices added into self-attention layers that require even fewer parameters, has also gained popularity [36]. Additionally, AdaptFormer [15], which integrates a scalable Adapter into the multi-layer perceptron (MLP) layer of the attention block specifically for vision transformers, has shown that adding just 2% extra parameters can surpass the performance of fully fine-tuned networks across various natural imaging benchmarks. In our study, we aim to investigate the effectiveness of applying PEFT for fine-tuning SAM within the medical imaging domain. Considering Adpater and LoRA are the most popular PEFT methods adopted in LLMs and vision foundation transformers [91], we choose to include them as well as the simplest Vanilla for our experiments. In summary, we select the following methods:

### v3

@@ Abstract
doi:
doi.org/10.59275/j.melba.2025-86a6keywords:
Machine Learning, Image Segmentation, Vision Foundation Model1 Introduction
@@
Medical image segmentation refers to identifying the region of interest, typically anatomical structures or tumors, in an image. This process is essential in the field of medical image analysis, contributing significantly to the effectiveness of computer-aided diagnosis (Al-Antari et al., 2018). The introduction of deep learning techniques has led to a variety of automated segmentation methods (Liu and Pang, 2020; Weng et al., 2023), further enhancing the precision and efficiency of analyzing medical images. Despite the progress made, it is still difficult to determine the most effective method for developing an optimal segmentation model, especially with the recent introduction of vision foundation models dedicated to segmentation. Therefore, in our study, we aim to explore and define the optimal strategies for developing automated segmentation algorithms specifically designed for medical images and with a focus on foundation models.
@@
In the era of deep learning, a common practice for automated medical image segmentation is to directly train UNet (Ronneberger et al., 2015) or its variants (Siddique et al., 2021; Wang et al., 2022) on a specific dataset and task. Recently, there has been growing interest in customizing vision foundation models for specific segmentation tasks (Zhang and Liu, 2023; Gu et al., 2024). Foundation models are typically transformer-based (Vaswani et al., 2017) neural networks pre-trained on extremely large datasets, allowing them to generalize to a wide range of tasks and datasets. The Segment Anything Model (SAM) (Kirillov et al., 2023) was proposed as the first large model designed specifically for segmentation tasks. SAM employs a prompt-based training method, aiming to generate segmentation masks for target objects in response to user-provided prompts, e.g., points or bounding boxes. It has demonstrated comparable or even superior zero-shot performance to fully supervised models in processing natural images. However, SAM performance is less impressive when applied directly to medical images (Mazurowski et al., 2023; Deng et al., 2023; Huang et al., 2024), and the requirement of prompts makes its direct usage with medical imaging segmentation tasks challenging.
@@
Removing the prompt requirement in SAM is relatively straightforward, as we can use “dummy” prompt embeddings as inputs during fine-tuning, a strategy verified by concurrent work (Zhang and Liu, 2023). However, the approaches proposed for adapting SAM to medical imaging tasks vary widely. Some works focused on adapting SAM’s decoder only (Yue et al., 2023), while others chose to modify the entire network (Zhang and Liu, 2023). Some works introduced an additional pre-training stage (Ma et al., 2024) using multiple medical images, while others usually do not. Though some prior survey work (Lee et al., 2024; Zhang and Jiao, 2023) have attempted to list different strategies and categorize them into various groups, questions remain regarding (1) which fine-tuned strategy is the most effective, (2) whether these models can achieve better performance than traditional UNets, and (3) whether using additional data is helpful.
@@
In this work, we aim to answer these questions by considering three common dataset availability scenarios in the medical imaging field: (1) only a single labeled dataset; (2) multiple labeled datasets for different tasks; and (3) multiple labeled and unlabeled datasets. The first scenario reflects the most common deep learning scenario of adapting SAM to a dataset of interest, and the latter two scenarios offer the opportunity to first incorporate broad medical domain knowledge into SAM using supervised or self-supervised training before adapting it to a specific dataset. Figure 1 summarizes the strategies and experiment setups we include. To avoid terminology confusion, we name these strategies as task-specific supervised training, task-expansive prompt-based learning, and task-agnostic self-supervised learning respectively.
@@
When considering the design choice during task-specific supervised training, our work aligns with several concurrent works (Lee et al., 2024; Zhang and Jiao, 2023; Zhang et al., 2024b). However, our work differs from theirs in that we conduct comprehensive experiments over medical imaging datasets across various modalities for a fair comparison, rather than simply compiling disparate results, and provide an organized and new catalog that could summarize existing approaches. The choice of each category is shown in Figure 2 and will be discussed in Section 2.2. Our category definition reflects the common design choices when building a segmentation algorithm from SAM, and we can categorize most existing works based on the proposed criteria, shown in Table 1.
@@
When incorporating extra medical knowledge into SAM, we explore two main setups of pre-training approaches: task-expansive prompt-based learning and task-agnostic self-supervised learning, depending on the availability of labels. Note that our objective is not to propose a novel pre-training strategy, as such strategies are often constrained by the availability of computation resources and datasets. Instead, our focus is on determining if additional pre-training with more available medical data is helpful and, if so, what the best combination of pre-training and task-specific fine-tuning is. Therefore, we adopt the strategies used during SAM’s development, and one of the most commonly applied pre-training strategies in this field. Namely, we use the same interactive training pipeline where prompts are simulated from the masks for task-expansive prompt-based learning; and masked autoencoding (MAE) (He et al., 2022) as the objective for task-agnostic self-supervised learning. We evaluate these approaches on the same 17 datasets.
@@
Our study is extended to two crucial scenarios when deploying a segmentation algorithm for annotation purposes: few-shot learning and interactive segmentation. The first scenario reflects the case of annotating a new dataset with only a limited number of annotations, while the latter examines the effectiveness of using tools to assist the annotation process. In both scenarios, we limit the evaluation datasets to MRI. When conducting few-shot learning, we limit the number of training samples to 5 and re-examine the effectiveness of setups 1-3 described in Figure 1. In the interactive segmentation scenario, we focus on task-specific supervised training with either boxes or points as input prompts.
@@
In conclusion, we systematically study different approaches to fine-tune SAM in the medical imaging field. Our paper investigates an important question: What is the best practice to develop an automated segmentation algorithm based on foundation models? We find the answers depend on the availability of the datasets and consider three scenarios: (1) only a single labeled dataset; (2) multiple labeled datasets (with different objects of interest); and (3) multiple labeled and unlabeled datasets. We propose task-specific supervised training, task-expansive prompt-based learning followed by task-specific supervised training, and task-agnostic self-supervised learning followed by task-specific supervised training as the solution to each scenario, respectively. By conducting systematic experiments on each strategy, we observe several useful guidelines and summarize them in Section 5.
@@ 2.1 SAM and removal of prompts
To begin with, we first briefly present an overview of the Segment Anything Model (SAM) (Kirillov et al., 2023) before introducing the fine-tuning strategies. SAM consists of three components: (1) image encoder, (2) prompt encoder, and (3) mask decoder, as illustrated in Figure 2 of the overview of SAM’s architecture. The image encoder, built on a robust Vision Transformer (ViT) framework, effectively converts a 2D image into a latent feature representation. Notably, in SAM’s original architecture, the image encoder offers flexibility in model size, with options including ViT-H(uge), ViT-L(arge), and ViT-B(ase) (Dosovitskiy et al., 2020). The prompt encoder is flexible, accommodating different prompt types such as point, box, or mask prompts. Depending on the input type, it generates either sparse embeddings for point/box prompts or dense embeddings for mask prompts. Lastly, the mask decoder is a transformer decoder using two-way cross-attention that integrates image and prompt embeddings to produce a multi-channel mask output, each channel representing varying confidence levels. SAM is trained on SA-1B, a dataset collected by the same team behind its development, which contains more than 1B masks from 11M natural images. During training, first, with equal probabilities, box and point prompts are generated. Point prompts are sampled uniformly from the ground truth mask, and box prompts are taken as the ground truth mask’s bounding box, with random noise added in each coordinate. The objective is the combination of focal loss (Ross and Dollár, 2017) and dice loss (Milletari et al., 2016).
@@ 2.1.1 Interactive segmentation to automated segmentation
where the “sparse_embeddings” and “dense_embeddings” serve as inputs to SAM’s mask decoder. Empirical evidence shows the effectiveness of this simple approach (Zhang and Liu, 2023). More advanced methods, including adding a detection network to automatically generate box prompts or a self-prompt module into SAM’s structure (Wu et al., 2023b; Gao et al., 2023; Shaharabany et al., 2023), are feasible but will introduce extra complexity into SAM’s structure, hampering our analysis on the effectiveness of fine-tuning strategies. Moreover, we believe our work offers a potentially effective strategy, i.e., feeding the generated prompts to the fine-tuned SAM instead of the original one, for this research direction.
@@ 2.2.1 Image Encoder Architecture
In this work, we select ViT-B and ViT-H from SAM’s provided checkpoints because they represent the most efficient and effective versions of SAM’s encoder, respectively. Additionally, we explore a more compact option: ViT-T(iny). This backbone is introduced in MobileSAM (Zhang et al., 2023a), where the authors distill the knowledge from SAM’s pre-trained ViT-H to ViT-T, a customized vision transformer with only of the network parameters, and achieve an inference speed that is times faster compared to the original SAM. The lightweight nature can widen the application of SAM, ease the computational requirements, and, more importantly, raise interesting research questions, such as whether distillation can maintain or even lead to the superior performance of SAM during fine-tuning.
@@ 2.2.2 Model Component
Some methods state the effectiveness of maintaining the encoder’s weights and only adapting the lightweight mask decoder of SAM’s architecture, as shown in Table 1. These approaches are grounded in the understanding that SAM’s encoder, trained on a comprehensive collection of natural images, is sufficiently powerful and knowledgeable in the general imaging domain, with zero-shot ability to perform feature extraction. Updating the lightweight decoder can enhance performance in specific applications. Also, it is suggested that fine-tuning the encoder of SAM would demand considerably higher computational resources. However, due to the significant appeal difference between natural images and medical images, it remains unclear if SAM’s encoder could effectively capture the distinctive features of medical images without any modifications to the image encoder. Therefore, we consider these two choices: mask decoder only and mask decoder plus image encoder.
@@
| Decoder + Vanilla | SAM for medical images? (Huang et al., 2024) | COSMOS 1050K | ViT-H / ViT-B | |
| Generalist Vision Foundation (Shi et al., 2023a) | Multiple | ViT-H | ||
| How to efficiently (Hu et al., 2023b) | ACDC (Bernard et al., 2018) | ViT-H | ||
| All-in-SA (Cui et al., 2023) | Monuseg (Kumar et al., 2019) | ViT-H | ||
| AdaptiveSAM (Paranjape et al., 2023) | Multiple | ViT-B | ||
| SkinSAM (Hu et al., 2023a) | HAM10000 (Tschandl et al., 2018) | ViT-B | ||
| DeSAM (Gao et al., 2023) | Multi-Prostate | ViT-H | ||
| SurgicalSAM (Yue et al., 2023) | EndoVis (Allan et al., 2020) | ViT-H | ||
| SAM3D (Bui et al., 2023) | Multiple | ViT-B | ||
| SAM Fewshot (Xie et al., 2024) | Multiple | N/A | ||
| Decoder + LoRA | BLO-SAM (Zhang et al., 2024a) | Multiple | ViT-B | |
| Encoder + Vanilla | AutoSAM (Shaharabany et al., 2023) | Multiple | ViT-H | |
| SAMUS (Lin et al., 2023) | Multiple | N/A | ||
| En/Decoder + Vanilla | Polyp-SAM (Li et al., 2023b) | Multi-Colonoscopy | ViT-L / ViT-B | |
| CellViT (Hörst et al., 2023) | Multi-nuclei | ViT-H | ||
| En/Decoder + LoRA | Customized SAM (Zhang and Liu, 2023) | Multi-Organ | ViT-H / ViT-B | |
| MA-SAM (Chen et al., 2023a) | BTVC (Fang and Yan, 2020) | ViT-H | ||
| Cheap Lunch (Feng et al., 2023) | BraTS 18 (Menze et al., 2014) | ViT-B | ||
| En/Decoder + Adapter | Medical SAM Adapter (Wu et al., 2023a) | BTVC (Fang and Yan, 2020) | ViT-H | |
| SAM Fails to Segment? (Chen et al., 2023b) | kvasir-SEG (Jha et al., 2020) | ViT-H | ||
| Learnable Ophthalmology SAM (Qiu et al., 2023) | Multiple | N/A | ||
| Auto-Prompting SAM (Li et al., 2023a) | Multiple | LightWeight ViT | ||
| SAM-Med2D (Cheng et al., 2023) | SAM-Med2D (Cheng et al., 2023) | ViT-B | ||
| 3DSAM-adapter (Gong et al., 2023) | Multi-CT | ViT-B | ||
| SAM-Path (Zhang et al., 2023b) | Multiple | LightWeight ViT | ||
| Self-Sampling Meta (Leng et al., 2024) | ABD-30 (Landman et al., 2015) | ViT-B | ||
| Cross-modality (Shi et al., 2023b) | MRI | ViT-B |
@@ 2.2.3 Vanilla or Parameter-efficient Fine-tuning Method
When conducting model fine-tuning or transfer learning, the easiest and most intuitive approach is to update all model parameters involved when feeding medical domain data, which is known as vanilla fine-tuning (Azizi et al., 2022).
Recently, parameter-efficient fine-tuning (PEFT) techniques have gained increasing attention. Compared with vanilla fine-tuning, PEFT keeps most of the network parameters fixed and only updates a small percentage, i.e. less than , of them. Such a strategy significantly reduces computational cost and is beneficial in preventing both catastrophic forgetting and overfitting (Dutt et al., 2023). The simplest PEFT techniques involve selecting and fine-tuning a subset of the parameters from the original pre-trained model. These selective methods concentrate on identifying the optimal parameters for fine-tuning, aiming to strike an ideal balance between parameter efficiency and overall model performance. For instance, BitFit (Zaken et al., 2021) proposed fine-tuning only the bias term in a transformer-based network, and Tinytl (Cai et al., 2020) shares a similar idea but focuses on CNN-based networks. Touvron et al. (Touvron et al., 2022) have demonstrated the effectiveness of selectively fine-tuning only the attention layers in Vision Transformers (ViTs), whereas alternative approaches suggest that exclusively optimizing the normalization layers (Basu et al., 2023).
Rather than updating certain selected layers, recent research (Li et al., 2022) indicates that incorporating extra modules into the large models appears to be more beneficial because it reconfigures models towards a downstream task rather than solely updating network parameters. Among these additional module methods, Adapters, which were originally introduced for learning multi-domain representations (Rebuffi et al., 2017, 2018), have emerged as a particularly prevalent choice. In addition to the conventional Adapter block (Houlsby et al., 2019), Low-Rank Adaptation (LoRA), known for its efficient low-rank matrices added into self-attention layers that require even fewer parameters, has also gained popularity (Hu et al., 2021). Additionally, AdaptFormer (Chen et al., 2022), which integrates a scalable Adapter into the multi-layer perceptron (MLP) layer of the attention block specifically for vision transformers, has shown that adding just 2% extra parameters can surpass the performance of fully fine-tuned networks across various natural imaging benchmarks. In our study, we aim to investigate the effectiveness of applying PEFT for fine-tuning SAM within the medical imaging domain. Considering that Adapter and LoRA are the most popular PEFT methods adopted in LLMs and vision foundation transformers (Xu et al., 2023), we choose to include them, as well as the simplest Vanilla, in our experiments. The implementation details of Adapter and LoRA are presented in Figure 2, Selection 3. Specifically, Adapter introduces two additional layers in each transformer block, and LoRA adds trainable pairs of rank decomposition matrices in parallel to existing blocks. Following (Hu et al., 2021), we set the rank of the additional matrices to 4, which provides a good balance between performance and computational efficiency. In summary, we select the following methods: